### v4

# These $\equalcontThese authors contributed equally to this work.$

Abstract
Decision-making is an essential attribute of any intelligent agent or group. Natural systems are known to converge to optimal strategies through at least two distinct mechanisms: collective decision-making via imitation of others, and individual trial-and-error. This paper establishes an equivalence between these two paradigms by drawing from the well-established collective decision-making model of nest-hunting in swarms of honey bees. We show that the emergent distributed cognition (sometimes referred to as the hive mind) arising from individual bees following simple, local imitation-based rules is that of a single online reinforcement learning (RL) agent interacting with many parallel environments. The update rule through which this macro-agent learns is a bandit algorithm that we coin Maynard-Cross Learning. Our analysis implies that a group of cognition-limited organisms can be equivalent to a more complex, reinforcement-enabled entity, substantiating the idea that group-level intelligence may explain how seemingly simple and blind individual behaviors are selected in nature.
From a biological perspective, this analysis suggests how such imitation strategies evolved: they constitute a scalable form of reinforcement learning at the group level, aligning with theories of kin and group selection. Beyond biology, the framework offers new tools for analyzing economic and social systems where individuals imitate successful strategies, effectively participating in a collective learning process. In swarm intelligence, our findings will inform the design of scalable collective systems in artificial domains, enabling RL-inspired mechanisms for coordination and adaptability at scale.
keywords:
Swarm Intelligence, Reinforcement Learning, Evolutionary Game Theory, Opinion Dynamics
1 Introduction
NN-parallel environments
RL update
“Hive mind”
Imitation
update
$n$-options
$n$-arms
Policy($\pi$)
Population $\mathcal{P}$
of NN-bees
$a_{B}\sim$$\mathcal{P}$
$a_{C}\sim$$\mathcal{P}$
$a_{A}\sim$$\mathcal{P}$
$a_{B}\sim$$\pi$
$a_{D}\sim$$\mathcal{P}$
Figure 1: The “hive mind” of a swarm of $N$ bees nest-hunting among $n$ options is a single $n$-armed bandit RL agent learning from NN environments in parallel.
Decision-making is the ability to choose the best action amongst all available options in a given scenario. Defining what a “best action” means requires ranking all possible outcomes [1]. Most often, the preference among outcomes is abstracted out as a single scalar signal [2] conveyed to the agent/individual [^1] To avoid confusion, we use “agent” in the context of RL and “individual” in the context of a population. by the environment as feedback for making a decision, which casts optimal decision-making into a maximization problem.
In nature, at least two distinct mechanisms allow single agents or groups of individuals to converge towards making optimal decisions. The first of these mechanisms adopts an individualistic approach, whereby the agent learns to make optimal decisions through trial and error. This paradigm is known as Reinforcement Learning (RL), where the agent learns a policy that maximizes the rewards it receives upon interacting with its environment [3]. Numerous studies in computational neuroscience have established links between learning processes happening in the brains of living beings and the formal framework of algorithmic RL [4, 5]. In this paper, we are specifically interested in the multi-armed bandit framework [3], where a single RL agent learns to make choices among different options (or “arms”) in an online fashion. Online RL agents learn by either interacting with a single environment (a setting recently referred to as the “streaming” setting [6, 7]), or, when possible, by collecting samples in parallel from multiple copies of the same environment simultaneously (the “parallel” setting [8]). To stabilize RL, the learning rate and the number of parallel environments are crucial in the streaming and parallel cases, respectively. Among the many learning algorithms designed to solve multi-armed bandits (Upper-Confidence-Bound [9], $\epsilon$-greedy [3], Gradient Bandit [10], etc.), we consider the Cross Learning (CL) [11] update rule (named after the economist John G. Cross [12] and not to be mixed with more recent overlapping nomenclature). CL is closely related to the Gradient Bandit algorithm, an RL algorithm directly based on the policy gradient theorem [3]. The policy gradient theorem is also the basis of many more advanced policy gradient (e.g., REINFORCE [13] and PPO [14]) and actor-critic (e.g., SAC [15]) RL algorithms.
The second mechanism takes a collective approach, where individuals mimic other individuals in the group to make decisions. Typically, this process is cast as a consensus problem, in which all individuals in a population have to agree on the best decision from a set of alternatives. This problem is commonly referred to as best-of-n collective decision-making (CDM). CDM arises in various domains, such as honey bee colonies [16, 17], human societies [18], and robot swarms [19, 20]. Imitation of success [21], from Evolutionary Game Theory (EGT), is a solution to CDM where individuals mimic randomly chosen neighbors based on how successful those neighbors’ decisions have been. Another solution to CDM is the weighted voter model [20, 22], which comes from modeling the nest-hunting behavior of honey bees [23]. During nest-hunting, honey bees are known to adopt the behavior represented in Fig. 1 (left-hand half): after scouting one of $n$ potential nesting areas, bees come back to the initial location of the swarm and perform a “waggle dance” [23] that describes the coordinates of the option they have explored. This dance is performed at a frequency that is proportional to the estimated quality of the explored area. Other bees go scout the area corresponding to the first dance they witness, and this process repeats until the colony reaches a quorum [24]. At this point, the entire swarm takes off and leaves for the winning site (note this is a simplified description of bees’ behavior, see [25, 24, 26, 27] for details). We refer to this distributed behavior as that of a “hive mind”, motivated by two factors: (1) biologists have already witnessed parallels between the emergent distributed cognition arising from individual-to-individual interactions and cognition in neuron-based complex brains of vertebrates [27], and (2): we formally show that the bee colony collectively learns as if it were one single RL entity.
In this paper, we highlight how imitation mechanisms in groups and individual trial-and-error mechanisms can be bridged via different variants of the Replicator Dynamic [28] (RD). In particular, we show that the model used by biologists to describe the nest-hunting behavior of honey bees amounts to a single-agent RL process at the swarm level, as illustrated in Fig. 1. To the best of our knowledge, this is the first work to formally establish this equivalence. Specifically, our contributions are:
  • We first remark that a large population of non-learning individuals following the imitation of success model can equivalently be seen as a single abstract RL agent following the Cross Learning update rule.
  • We then show that the weighted voter model used to model the behavior of individual honey bees during nest-hunting also aggregates to a single-agent RL algorithm at the macro-organism level, that we coin Maynard-Cross Learning.
  • We extend our theoretical analysis with further results from simulation.
The first two contributions, which detail how social imitation models aggregate into a single reinforcement learning agent, provide a formal description of diverse instances of collective intelligence.
2 Background
2.1 Multi-armed bandits and Cross Learning (J.G. Cross)
Multi-armed bandits (Fig. 1, right-hand half) are the simplest type of environment encountered in RL literature. They consist of a discrete set of available actions, called “arms”, among which the agent has to find the most rewarding. In the $n$-armed bandits considered in this paper, pulling an arm $a\in\{1,\dots,n\}$ returns a real-valued reward $r_{a}\in[0,1]$ sampled from a hidden distribution $r(a)$. The objective for an RL agent playing a multi-armed bandit is to learn a policy, denoted by the probability vector $\pi=(\pi_{1},\dots.\pi_{n})$, that maximizes the rewards obtained upon pulling the arms. Different optimization strategies exist to find such policies, one of the oldest being Cross Learning [11]:
Definition 1.
Let $k$ be an action and $r_{k}$ a corresponding reward sample ($r_{k}\sim r(k)$). Let $\pi_{a}$ denote the $a^{\text{th}}$ component of $\pi$. Cross Learning (CL) updates the policy $\pi$ as: $\forall a,\pi_{a}\leftarrow\pi_{a}+r_{k}\times\begin{cases}1-\pi_{a}&\text{if}\ a=k\\ -\pi_{a}&\text{otherwise}\end{cases}$ (1)
For convenience, when sampling reward $r_{k}$ from action $k$, we denote the expected policy update on action $a$’s probability $\pi_{a}$ as:
$d\pi_{a}(k)=\mathbb{E}_{r_{k}\sim r(k)}[r_{k}]\times\begin{cases}1-\pi_{a}&\text{if}\ a=k\\ -\pi_{a}&\text{otherwise}\end{cases}$ (2)
In CL, every reward $r_{k}$ sampled by applying an action $k$ directly affects the probabilities of all actions. As noted earlier, CL is close to the Gradient Bandit algorithm, which performs a similar update at the parameter level (called “preferences”) of a parametric policy rather than directly updating the probability vector.
2.2 Evolutionary Game Theory
Evolutionary Game Theory (EGT) is the study of population games [21]. A population $\mathcal{P}$ is made of a large number of individuals, where any individual $i$ is associated with a type, denoted by $T_{i}\in\{1,\dots,n\}$. The population vector $\pi=(\pi_{1},\dots,\pi_{n})$ represents the fraction of individuals in each type ($\sum_{i}\pi_{i}=1$). Individuals are repeatedly paired at random to play a game, each receiving a separate payoff. Individuals adapt their type based on their payoff according to a revision protocol.
Remark 1.
Population-policy equivalence. For the argument of our paper, it is interesting to interchangeably define $\pi=(\pi_{1},\dots,\pi_{n})$ both as a multi-armed bandit RL policy and as a population vector [12]. This is possible because in both cases the vector $\pi$ is constrained to the probability simplex. Further, note that uniformly sampling an individual of type $a$ from the population $\mathcal{P}$ (represented by the population vector $\pi$) is equivalent to sampling an action $a$ from the policy $\pi$.
2.2.1 Imitation of success and the Taylor Replicator Dynamic
In evolutionary dynamics, an important revision protocol is imitation of success [21]:
Definition 2.
In the “imitation of success” revision model $R_{\text{success}}$, any individual $i\in\mathcal{P}$ of type $T_{i}=a$ executes the following process: $i$ samples a random individual $j\sim\mathcal{U}(\mathcal{P})$ to imitate. Let $T_{j}$ be $b$. Both individuals $i$ and $j$ play a 2-player game in which they receive payoffs $r_{a}$ and $r_{b}$ respectively ($0\leq r_{a,b}\leq 1$). Each payoff depends on the types of both individuals. $i$ switches from type $a$ to type $b$ with probability $r_{b}$.
One can easily see why this rule is called “imitation of success”: $i$ imitates $j$ based on $j$’s payoff. Imitation of success is typically used to model replication of the fittest in evolution for biology, or certain human behaviors in economics [29]. When aggregated across the population, this revision model yields a famous evolutionary dynamic known as the Taylor Replicator Dynamic [30, 28] (TRD) (see LABEL:lem:PG):
$\dot{\pi}_{a}=\pi_{a}(q^{\pi}_{a}-v^{\pi}),$ (3)
where $\dot{\pi}_{a}$ is the derivative of the $a$-th component of the population vector, $q^{\pi}_{a}:=\mathbb{E}[r_{a}]$ is the expected payoff of the type $a$ against the current population, and $v^{\pi}:=\sum_{b}\pi_{b}\mathbb{E}[r_{b}]$ is the current average payoff of the entire population.
2.2.2 Weighted Voter Model and the Maynard-Smith Replicator Dynamic
The weighted voter model [22] instead typically models the nest-hunting behavior of honey bees described in Section 1 [19, 23]:
Definition 3.
In the “weighted voter” revision model $R_{\text{wvoter}}$, any bee $i\in\mathcal{P}$ of type $T_{i}=a$ (where $a$ corresponds to a nest-site option) executes the following process: $i$ estimates the quality of its current type $r_{a}\sim r(a)$, where $0\leq r_{a}\leq 1$. After obtaining $r_{a}$, $i$ locally broadcasts its type at a frequency proportional to $r_{a}$. $i$ switches its type to the first type $b$ that it perceives from its neighborhood. Assuming all individuals are well mixed in the population [31], the corresponding expected probability of $i$ switching to type $b$ is the proportion of votes cast for $b$ within its neighborhood: $P^{(i)}_{\text{neighborhood}}(b\leftarrow a)=\frac{N^{(i)}_{b}\mathbb{E}[r_{b}]}{\sum_{l}N^{(i)}_{l}\mathbb{E}[r_{l}]}$ where $N^{(i)}_{k}$ is the number of individuals of type $k$ in the neighborhood of $i$.
Note that, in this model, honey bees do not need to directly observe the quality estimate of other scout bees, but only their type. In Section 3.2, we show that the weighted voter revision model aggregates to a variant of the TRD, called the Maynard-Smith Replicator Dynamic [32] (MRD):
$\dot{\pi}_{a}=\frac{\pi_{a}}{v^{\pi}}(q^{\pi}_{a}-v^{\pi})$ (4)
3 Methodology
In Section 3.1, we remark that previous results from the literature yield an interesting insight binding $R_{\text{success}}$ and reinforcement learning together: a large population following the “imitation of success” revision model can equivalently be considered as a single RL agent. While this is a direct consequence of previously known results, we could not find this insight formulated in the literature and thus we took the liberty to formalize it as 1. Then, in Section 3.2, we prove that the weighted voter revision model also aggregates to an RL algorithm. In other words, our analysis indicates that, at least in the well-studied case of nest-hunting, a swarm of honeybees collectively acts as a single RL entity. We formalize this result as 2.
3.1 Imitation of Success and Cross Learning
Evolutionary Game Theorists have long been interested in the “imitation of success” revision protocol, as it models replication of the fittest in the evolutionary setting. In this literature, it is famously known that a population of individuals following $R_{\text{success}}$ aggregates to the Taylor Replicator Dynamic (see for instance [21, 28] and the proof in the Appendix):
Lemma 1.
lem:PG An infinite population of individuals adopting $R_{\text{success}}$ follows the TRD:
$d\pi_{a}=\pi_{a}(q^{\pi}_{a}-v^{\pi}),$ (5)
where $\pi_{a}$ is the proportion of type $a$ in the population, $q^{\pi}_{a}$ is the expected payoff (also called “fitness”) of type $a$ against the population, and $v^{\pi}$ is the average fitness of the population.
Our choice of notation in LABEL:lem:PG is reminiscent of the RL literature and may seem unusual from an EGT perspective. In fact, this choice is motivated by another, lesser-known result from [33, 12], who showed that the Cross Learning RL algorithm performs updates that are also similar to the TRD (proof in Appendix):
Lemma 2.
lem:RL In expectation, an RL agent learning via the CL update rule follows:
$\mathbb{E}[d\pi_{a}]=\pi_{a}(q^{\pi}_{a}-v^{\pi}),$ (6)
where $q^{\pi}_{a}$ is the action-value of $a$, and $v^{\pi}$ is the value of policy $\pi$.
With this notation, it is straightforward to combine LABEL:lem:PG and LABEL:lem:RL, provided a sensible duality exists between the identical terms. Remember how Remark 1 describes a duality between the population vector of LABEL:lem:PG and the policy of LABEL:lem:RL, $\pi$. In LABEL:lem:PG, the fitness $q^{\pi}_{a}$ is the payoff that an individual of type $a$ can expect on average when encountering a random individual from the population $\pi$. In LABEL:lem:RL, the dual of this individual of type $a$ is an action $a$ sampled from the policy $\pi$, whose expected reward is the action-value $q^{\pi}_{a}$. Similarly, the dual of the population fitness $v^{\pi}$ is the policy-value. In other words, under $R_{\text{success}}$, individuals sampled from the population $\pi$ can equivalently be regarded as action samples from an RL macro-agent, whose rewards are the individuals’ payoffs. The agent maximizes the average payoff of the group via Exact Cross Learning (we call “exact” the RL algorithm that directly applies expected updates instead of updates computed from sample estimates):
Proposition 1.
An infinite population of individuals following $R_{\text{success}}$ can equivalently be seen as an RL agent following Exact Cross Learning, i.e., $d^{\text{success}}\pi_{a}=\mathbb{E}[d^{\text{CL}}\pi_{a}]\;,$ (7) where $\pi$ is both a population vector and a vector of action-probabilities, $d^{\text{success}}\pi$ is the single-step change in the population vector $\pi$ under the “imitation of success” revision model , and $d^{\text{CL}}\pi$ is an update performed by CL on the policy $\pi$.
Proof.
Direct consequence of LABEL:lem:PG and LABEL:lem:RL. ∎
Intuitively, this macroscopic RL process can appear as a mere by-product of individuals in the population imitating others to optimize for their own success. But let us now turn our attention to organisms that may not even have a notion of individual success to optimize: honey bees.
3.2 Weighted Voters and Maynard-Cross Learning
We now consider a large population of $N\gg 1$ honey bees with average local neighborhood size $M\gg 1$, seeking agreement on which nesting site to select by applying $R_{\text{wvoter}}$. We show that, although $R_{\text{wvoter}}$ is a blind imitation protocol where individual bees simply imitate the first type they encounter, a swarm of bees following $R_{\text{wvoter}}$ aggregates to an RL agent at the macro-organism level. In the $R_{\text{wvoter}}$ model, individual bees are blind imitators, but active promoters. Each bee $i$ of type $T_{i}=k$ and payoff sample $r^{(i)}_{k}\sim r(k)$ has a tangible stochastic influence on the local expected inflow of other bees that it rallies to its own type $k$ within its local neighborhood $\mathcal{N}_{i}$:
$P^{(i)}_{\text{neighborhood}}(k\leftarrow\cdot)=\frac{r^{(i)}_{k}}{\sum\limits_{j\in\mathcal{N}_{i}}r^{(j)}},$ (8)
where $r^{(i)}_{k}$ represents $i$’s broadcasting frequency, and $\sum_{j\in\mathcal{N}_{i}}r^{(j)}$ represents the total broadcasting frequency of $i$’s local neighborhood. The expected outflow attributable to $i$ on the entire swarm from any type $a\neq k$ to the type $k$ is thus:
$P^{(i)}(k\leftarrow a)=\frac{N^{(i)}_{a}}{N}\frac{r^{(i)}_{k}}{\sum\limits_{j\in\mathcal{N}_{i}}r^{(j)}},$ (9)
where $N^{(i)}_{a}$ is the number of type-$a$ bees within $i$’s neighborhood. For simplicity, we assume that $M$ (and thus also $N^{(i)}$) is large. Assuming that all individuals are well mixed in the population [31], it follows that $\frac{N^{(i)}_{a}}{M}=\pi_{a}$:
$P^{(i)}(k\leftarrow a)$ $=\frac{M\pi_{a}}{N}\frac{r^{(i)}_{k}}{\sum\limits_{j\in\mathcal{N}_{i}}r^{(j)}}$
$=\frac{\pi_{a}}{N}\frac{r_{k}}{\frac{1}{M}\sum\limits_{j\in\mathcal{N}_{i}}r^{(j)}}$
$=\alpha\frac{r_{k}}{v^{\pi}}\pi_{a},$ (10)
where $v^{\pi}$ is the average population payoff and $\alpha:=\frac{1}{N}$. Summing over all types except $k$ (whose outflow to $k$ is 0), we obtain the total inflow into type $k$ attributable to $i$:
$\sum_{a\neq k}P^{(i)}(k\leftarrow a)$ $=\sum_{a\neq k}\alpha\frac{r_{k}}{v^{\pi}}\pi_{a}$
$=\alpha\frac{r_{k}}{v^{\pi}}(1-\pi_{k})\;.$ (11)
Given a bee $i$ of type $k$, Eq. 10 describes the outflow that its waggle dance induces within the swarm from any type $a$ to the type $k$, while Eq. 11 describes the corresponding inflow into type $k$ from all types. In other words, the bee $i$ can be seen as an action-reward sample whose influence on the population vector $\pi$ is described by an RL update rule that we coin $\alpha$-Maynard-Cross Learning ($\alpha$-MCL):
Definition 4.
Let $k$ be an action and $r_{k}\sim r(k)$ a corresponding reward sample. MCL updates the policy $\pi$ as: $\forall a,\pi_{a}\leftarrow\pi_{a}+\alpha\frac{r_{k}}{v^{\pi}}\begin{cases}1-\pi_{a}&\text{if}\ a=k\\ -\pi_{a}&\text{otherwise}\end{cases}$ (12) where $v^{\pi}$ is the current value of policy $\pi$.
Here, the meaning of $\alpha$ is of peculiar interest. In Eq. 12, $\alpha$ looks very much like a learning rate, i.e., a hyperparameter that RL practitioners typically set to a small value in order to downsize the amplitude of individual policy updates. When dealing with on-policy RL algorithms, computer scientists often evaluate a policy in several parallel copies of the same environment [34], average the corresponding updates into a single policy update, and downsize this update via a small learning rate. The common rule of thumb is that the learning rate can be larger when there are enough parallel environments. But in Definition 4, $\alpha$ instead has a clear population-based meaning: it is the inverse of the number of bees acting in parallel in the swarm, as we are describing the influence of one single bee on the population. Crucially, this influence is that of an RL agent, but the individual bee itself is in no way an RL agent: it only blindly imitates its peers! Instead, everything happens as if nature had evolved individual bees’ behavior such that it aggregates to an RL agent at the macro-organism level: the “hive mind”, whose rewards are the quality estimates of nest site samples. Summing the influence of all individual bees yields:
$d\pi_{a}$ $=\sum_{i=0}^{N}\frac{r^{(i)}}{Nv^{\pi}}\begin{cases}1-\pi_{a}&\text{if}\ a=k\\ -\pi_{a}&\text{otherwise}\end{cases}$
$=\sum_{k}\frac{N_{k}q^{\pi}_{k}}{Nv^{\pi}}\begin{cases}1-\pi_{a}&\text{if}\ a=k\\ -\pi_{a}&\text{otherwise}\end{cases}$
$=\frac{1}{Nv^{\pi}}(N_{a}q^{\pi}_{a}(1-\pi_{a})-\sum_{k\neq a}N_{k}q^{\pi}_{k}\pi_{a})$
$=\frac{1}{Nv^{\pi}}(N_{a}q^{\pi}_{a}-\sum_{k}N_{k}q^{\pi}_{k}\pi_{a})$
$=\frac{1}{v^{\pi}}(\pi_{a}q^{\pi}_{a}-\pi_{a}\sum_{k}\pi_{k}q^{\pi}_{k})$
$=\frac{\pi_{a}}{v^{\pi}}(q^{\pi_{a}}-v^{\pi})$ (13)
which is the Maynard-Smith Replicator Dynamic. Similar to 1, we can further show that this update is the expected update performed by the 1-MCL algorithm, that we simply call Maynard-Cross Learning (MCL) for conciseness:
$\mathbb{E}[d^{\text{MCL}}\pi_{a}]$ $=\sum_{k=1}^{n}\pi_{k}.d^{\text{MCL}}\pi_{a}(k)$
$=\pi_{a}.d^{\text{MCL}}\pi_{a}(a)+\sum_{k\neq a}\pi_{k}.d^{\text{MCL}}\pi_{a}(k)$
$=\pi_{a}\frac{\mathbb{E}[r_{a}]}{v^{\pi}}(1-\pi_{a})+\sum_{k\neq a}\pi_{k}\frac{\mathbb{E}[r_{k}]}{v^{\pi}}(-\pi_{a})$
$=\frac{\pi_{a}}{v^{\pi}}\Bigr[\mathbb{E}[r_{a}]-\pi_{a}\mathbb{E}[r_{a}]-\sum_{k\neq a}\pi_{k}\mathbb{E}[r_{k}]\Bigr]$
$=\frac{\pi_{a}}{v^{\pi}}\Bigr[\mathbb{E}[r_{a}]-\sum_{k}\pi_{k}\mathbb{E}[r_{k}]\Bigr]$
$=\frac{\pi_{a}}{v^{\pi}}(q^{\pi}_{a}-v^{\pi})$ (14)
Thus we can write the RL update of the “hive mind” in a swarm of bees following the $R_{wvoter}$ model:
Proposition 2.
An infinite population of individuals following $R_{\text{wvoter}}$ can equivalently be seen as an RL agent following Exact Maynard-Cross Learning, i.e., $d^{\text{wvoter}}\pi_{a}=\mathbb{E}[d^{\text{MCL}}\pi_{a}],$ (15) where $\pi$ is both a population vector and a vector of action-probabilities, $d^{\text{wvoter}}\pi$ is the single-step change in the population vector $\pi$ under the weighted voter revision model, and $d^{\text{MCL}}\pi$ is the update performed by MCL on the policy $\pi$.
Proof.
Direct consequence of Eqs. 13 and 14. ∎
4 Final remarks and simulations
(a) $\alpha$-MCL for varying learning rates
(b) Varying swarm sizes NN
(c) Varying neighborhood sizes for NN=500
(d) Variants of $R_{\text{success}}$ & $R_{\text{wvoter}}$ in NN=1000
Figure 2: Simulations: (a) Swarms of bees reach consensus more rapidly when following $R_{\text{wvoter}}$ collectively than when individual bees learn via $\alpha$-MCL. (b,c) Varying swarm and neighborhood sizes show that the theoretical predictions hold under practical constraints. (d) Simple variants of $R_{\text{success}}$ and $R_{\text{wvoter}}$ can surpass $R_{\text{wvoter}}$ in this environment, raising the question of why evolution favored $R_{\text{wvoter}}$ over alternative decision-making strategies.
4.1 A single RL agent in many parallel environments
Researchers have already informally witnessed collective intelligence [35, 36] described in Section 3, often referred to as “emergent” [37] to describe situations where simple, seemingly blind individual behaviors aggregate to a complex and coherent behavior of the group. 2 shows that, when the individual bees’ behavior is modeled as $R_{\text{wvoter}}$, this “emergent” collective intelligence is MCL, a multi-armed bandit Reinforcement Learning algorithm. But the implications of 2 go even further: not only a swarm of bees is a single MCL agent, it is one that is quite efficient at what it does, because each bee is a parallel action sample of its policy. In other words, each bee $i$ can be seen as an action sample $a^{(i)}\sim\pi$ tested against a parallel copy $E^{(i)}$ of the environment $E$, as illustrated in Fig. 1. Fig. 2(a) shows that this converges much faster than if individual bees were themselves to learn via $\alpha$-MCL instead of just following $R_{\text{wvoter}}$. If the MCL algorithm were of any use outside of biology [^2] Most likely not: more advanced algorithms exist to solve multi-armed bandits (such as UCB [9])., say if a computer scientist wanted to train a robot via iterative $\alpha$-MCL, they would need a small $\alpha$ for the algorithm to converge to the optimal option. Or, if feasible, they could afford using a larger $\alpha$ by evaluating the policy in several parallel simulations (see LABEL:app:streaming_rl and LABEL:app:parallel_rl). Swarms of honey bees naturally learn via parallel RL in as many environments as there are scouts in the swarm. Finally, 1 implies that a similar discussion holds for the “imitation of success” revision protocol, which encompasses certain human behaviors, and even natural selection itself.
4.2 Swarm size
In Section 4.1, we have discussed how being made of a number of individuals enables the “hive mind” to evaluate actions sampled from $\pi$ in a parallel fashion. Furthermore, our analysis assumed that this number was large (which enabled us to compute flows in Section 3). Yet, in practice there is a tradeoff here. Actual honey bee swarms are known to use only a small proportion of their bees as scouts, while the large majority of the swarm remains quiescent during nest-hunting [38, 39], which is surmised to be in the interest of energy-efficiency. In a typical swarm of 10,000 bees, roughly 200 to 500 scouts are active during nest-site selection. To analyze how the number of scouts $N$ affects the task-performance of the MCL macro-agent, we conducted simplified simulations where swarms of varying sizes had to choose the best amongst 10 options. Option qualities were spread between 0 and 1, while the noise in bees’ estimates was modeled with a uniform perturbation of amplitude $0.2$. Fig. 2(b) reports the percentage of populations that converged to the optimal choice for different values of $N$, across 1000 seeds (each individual seed converged to a homogeneous choice). In these simulations, $N=500$ is enough to closely follow the MRD and converge to the optimal option, whereas a number of scout bees that is too small often yields convergence to a sub-optimal nest site option. However, our choice of simulation parameters is largely arbitrary: Fig. 2(b) should be interpreted qualitatively rather than quantitatively.
4.3 Neighborhood size
In Section 3, we have assumed that the local neighborhood size $M$ was large. This assumption was useful to analytically derive the macro-agent RL update, because it meant that local neighborhoods (i.e., other scout bees that a scout may randomly walk into within its vicinity) could be considered well-mixed. Arguably, actual neighborhood sizes are not that large in the real world. Therefore, to complete our theoretical analysis, we performed $R_{\text{wvoter}}$ simulations with varying neighborhood sizes. Fig. 2(c) shows that even a small $M\geq 5$ yields a macro-agent algorithm that closely follows MCL. With $M=1$, each scout copies an arbitrary neighbor, which yields no macro-dynamic.
4.4 The collective power of promoting
The $R_{\text{wvoter}}$ revision protocol adopted by honey bees has surprising advantages over imitation of success. In LABEL:sup:methods, we show that $R_{\text{wvoter}}$ (MCL) generally converges faster than $R_{\text{success}}$ (CL), except for small population sizes. This is in addition to the simplicity of $R_{\text{wvoter}}$, which only requires bees to blindly mimic their peers.
4.5 Could bees do more?
In Section 3, we showed that the bees employing $R_{\text{wvoter}}$ can be seen as a single online RL agent. However, a natural question arises: is $R_{\text{wvoter}}$ the optimal collective-decision making strategy? To answer this question, we investigate three simple variants of the imitation of success model and the weighted voter model (more details in LABEL:app:pop_details):
  • Deterministic Imitation of Success: Unlike the standard imitation of success model (which switches stochastically based on partner’s success), this variant deterministically adopts the partner’s type when it has more success.
  • Deterministic Imitation of Success with Weighted Voter Rule: This variant combines the weighted voter model’s success-weighted neighbor sampling with the deterministic imitation of success switch based on comparative success.
  • Stochastic Imitation of Success with Weighted Voter Rule: This variant differs from Deterministic Imitation of Success with Weighted Voter Rule in that it uses classic stochastic imitation of success instead of our deterministic version as the switching rule.
Fig. 2(d) shows that, at least in this simplified environment, all these variants converge to the optimal decision faster than $R_{\text{wvoter}}$ or MRD. This simple experiment suggests that there may be better collective decision-making strategies than $R_{\text{wvoter}}$. In particular, we see that if the bees could perform comparisons with their individual quality estimates before blindly imitating others from the neighborhood, the convergence speed could be faster than MRD. This insight raises an important question: why did bees evolve to use $R_{\text{wvoter}}$ [19, 23] over other collective decision-making strategies?
We speculate that this behaviour relates to cognitive constraints and the fundamental challenge of quality comparison. The weighted voter model requires remarkably minimal individual capabilities: bees need only estimate the quality of their own option and observe the types (not quality estimates) of neighbors, broadcasting their preference at a frequency proportional to quality. $R_{wvoter}$ requires no comparison operations between quality values and no direct communication of numerical quality estimates, capabilities that all faster-converging variants demand. In contrast, deterministic switching rules require bees to not only observe others’ quality estimates and compare them to their own, but also presume an objective or at least mutually calibrated scale for evaluating nest site quality. This is non-trivial: individual bees may experience different environmental conditions during their assessments, have different sensory acuity, or weight various quality features differently. $R_{wvoter}$ may elegantly sidestep this calibration problem by converting subjective quality estimates into broadcast frequencies — a transformation that preserves relative preferences while eliminating the need for interpersonal comparison. The weighted voter model may thus represent an evolutionary solution that achieves collective reinforcement learning with the absolute minimum individual cognitive machinery while avoiding the difficult problem of quality scale calibration, trading some convergence speed for dramatic simplification of individual processing and robustness to individual variation.
This hypothesis aligns with the broader pattern in social insects where complex collective behaviors emerge from remarkably simple individual rules, suggesting that evolutionary pressures favor cognitive parsimony at the individual level when collective intelligence can compensate at the group level, a strategy that may be driven in part by the substantial metabolic costs of neural tissue, which in mammals requires nearly an order of magnitude more energy per unit weight than other somatic tissues [40, 41]. We leave this exploration open for future work.
4.6 Broader significance
  • Biology. 2 shows that seemingly blind, imitation-based bee interactions aggregate to a form of collective intelligence, which is that of a “hive mind” learning through reinforcement. This is interesting from a neuroscience perspective, since both imitation and reinforcement have long been known to govern learning at the individual level [42, 4]. What our paper shows about honey bees is that their puzzling imitative behavior can in fact be understood as reinforcement from the swarm-agent perspective, where this behavior equates to massively parallel MCL. Why this imitative behavior has evolved in nature may thus appear clearer: swarms of honey bees are an example of reinforcement strategy having evolved at the group level. This analysis agrees with the commonly accepted theory of group/kin selection to explain the behavior of honey bees [43]. Building on the framework established in 1, parallel findings in other phylogenetically distant biological species underscore the generality of the proposed population–policy equivalences of this work. Notably, pheromone-guided C. elegans foraging exhibits the same mathematical structure as cross-learning [44], suggesting that stigmergic coordination may also be modeled as a form of collective reinforcement learning.
  • Economy and Social Dynamics. Evolutionary Game Theory goes beyond modeling biology: it extends to many types of population dynamics, most notably in economics [33, 29]. When thinking of $R_{\text{success}}$ as a model of certain decision-making strategies, 1 yields that imitating others’ success makes us part of a macroscopic RL process that hugely benefits the whole group, by mutualizing our speed of convergence to optimal strategies (see Section 4.1). This opens an avenue for modeling the group-level impacts of economic actors’ imitative behaviors as forms of collective reinforcement learning.
  • Algorithmic Swarm Intelligence. The field of Swarm Intelligence takes inspiration from complex emergent behaviors arising from a collective of natural entities following simple, local, and decentralized rules [35], to engineer algorithms that leverage collective principles like coordination, cooperation, and communication to tackle various problems (including CDM) across multiple domains such as swarm robotics [45, 46] and optimization [47]. Propositions 1 and 2 provide theoretical grounding for the phenomenon commonly referred to as “emergence” in this field [37]. This bridge also opens up the possibility of porting ideas from Reinforcement Learning to derive Swarm Intelligence local rules.
  • Algorithmic Reinforcement Learning. Similar to how we abstracted a population of $R_{wvoter}$ individuals as a single MCL agent, other revision models (majority rule [19] and cross-inhibition [48]) can be used to abstract dynamic groups of entities as simple, single RL agents. In massively multi-agent real-world settings (e.g., finance), this may greatly help model the non-stationarity of environment dynamics.
4.7 Conclusion
We have demonstrated that imitation-based collective behavior in large populations can be mathematically equivalent to reinforcement learning, providing a unifying framework that reinterprets swarm intelligence, social dynamics, and evolutionary processes as emergent forms of collective learning. This result provides a formal basis for understanding how simple, local imitation behaviors, often regarded as non-cognitive, can give rise to group-level intelligence and adaptive learning at scale. In particular, 2 demonstrates that the “hive mind” observed in swarms of bees can indeed be viewed as a single coherent agent learning via reinforcement, giving rise to a new bandit algorithm that we coin Maynard-Cross Learning.